The maximum efficiency of a steam turbine cycle with a steam temperature of 556oC and a cooling water temperature of 27oC is (give your answer in %) ______________.
Answer #1

T(hot) = 556 + 273 = 829K

T(cold) = 27 + 273 = 300K

Efficiency (%) = [T(hot) - T(cold)] x 100 / T(hot)

Efficiency (%) = [829 - 300] x 100 / 829

Efficiency (%) = 63.8%